OpenStudy (anonymous):

FAN AND MEDAL PLEASE HELP A circular dish has a diameter of 9 inches. What is the area of the dish? Use the formula a=pi r^2 A. 12.56pi in^2 B. 16pi in^2 C. 20.25pi in^2 D. 81 pi in^2 Please explain work step by step. Fan and medal

OpenStudy (anonymous):

The Area is given by teh formula: \(A = \pi r^2\)

OpenStudy (anonymous):

just plug r = 9 inches and find A

OpenStudy (anonymous):

But the diameter is 9 inches. It doesn't say the radius is 9 inches. Are they both 9 in this case?

OpenStudy (anonymous):

@M4thM1nd

OpenStudy (anonymous):

Ohh sorry, i thought i read radius, your are right, \(2r = 9\), so \(r = 9/2 = 4.5\)

OpenStudy (anonymous):

the radius is 4.5 right?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I just plug it in? if I just plug it in ill just get a number. I need the answer in terms of pi like in the multiple choice I put in the question

OpenStudy (anonymous):

i think it is c. 20.25pi

OpenStudy (anonymous):

Thats correct!